(a) Find fx(x,y), fy(x,y), fx(1,3), and fy(-2,4) for the given function. If

𝑧 = 𝑓(π‘₯, 𝑦) = 3π‘₯ΰ¬· 𝑦ଢ βˆ’ π‘₯ΰ¬Ά 𝑦ଷ + 4π‘₯ + 9

(b) A firm estimates that it can sell Q units of its product with an advertising expenditure of x thousand dollars where

𝑄 = 𝑄(π‘₯) = βˆ’π‘₯ΰ¬Ά + 600π‘₯ + 25

i) Over what level of advertising expenditure is the number of units of product sold increasing?

ii) Over what level of advertising expenditure is the number of units of product sold decreasing?Β 

1
Expert's answer
2021-06-16T13:16:08-0400

a)

"z=3x^3y^2-x^2y^3+4x+9"

"f_x=9x^2y^2-2xy^3+4"

"f_x(1,3)=9\cdot9-2\cdot27+4=31"

"f_y=6x^3y-3x^2y^2"

"f_y(-2,4)=-6\cdot8\cdot4-3\cdot4\cdot16=-384"


b)

"Q(x)=-x^2+600x+25"

"Q'(x)=-2x+600=0"

"x=300"


i)

"Q'(x)>0" for "0\le x<300" : the number of units of product sold increasing


ii)

"Q'(x)<0" for "x>300" : the number of units of product sold decreasing